jill is twice as old as his brother. the difference of their age is 45. how old is jill?
Answer by addingup(3677)

First, I think Jill is she and not he?
Second, you need to look at the problem and correct your numbers. I don't think th 45 is correct.
:
let's call Jill x and brother y:
x = 2y
x-y = 45 substitute for x
2y-y = 45
1y = 45
y = 45 the broher is 45 and Jill is 45*2= 90